OpenStudy (anonymous):

p(x)=x^3(x+2)(x-5)^2

OpenStudy (anonymous):

find the multiplicity and zeros of each

OpenStudy (anonymous):

it is the exponent

OpenStudy (anonymous):

\[x^3(x+2)(x-5)^2=0\] one factor is \(x^3\) which is zero if \(x\) is the "multiplicity " of that zero is three, because of the cube

OpenStudy (anonymous):

what do you do with the other exponent?

OpenStudy (anonymous):

so we would say "0 is a zero of multiplicity 3" now \(x+2=0\) if \(x=-2\) so \(-2\) is also a zero

OpenStudy (anonymous):

the multiplicity of that zero is one, because the factor is just \((x+2)\) you would say "\(-2\) is a zero of multiplicity one"

OpenStudy (anonymous):

how about \((x-5)^2\) ?

OpenStudy (anonymous):

yeah whats that all about

OpenStudy (anonymous):

lol solve \(x-5=0\) what do you get?

OpenStudy (anonymous):

is there any synthetic division/long division required in this problem?

OpenStudy (anonymous):

so the other zero is 5?

OpenStudy (anonymous):

no you do this problem only with your eyeballs no computation necessary at al yes, the other zero is 5

OpenStudy (anonymous):

what is the "multiplicity" of that zero?

OpenStudy (anonymous):

5? :/

OpenStudy (anonymous):

wait! 1!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the zero is 5 for the multiplicity recall that the factor is \((x-5)^2\)

OpenStudy (anonymous):

\[\large (x-\color{red}{5})^{\color{blue}2^{\text{ multiplcity is here}}}\]

OpenStudy (anonymous):

oh so the multiplicity is 2?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

see, you just look at it, it is displayed for you nothing at all to compute

OpenStudy (anonymous):

this is so sketchy. my teacher threw this in our homework as a curve ball because in the examples in class we had to divide and then factor out all this stuff and do the p/q thing to find the zeros

OpenStudy (anonymous):

yeah, but this problem starts at the end it is already in factored form, so there is nothing to do

OpenStudy (anonymous):

i.e. the factoring has already been done for you

OpenStudy (anonymous):

awesome that makes sense!

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

now i have to draw a graph

OpenStudy (anonymous):

ok it goes like this you have zeros at \(-2,0,5\) that is where the graph hits the \(x\) axis

OpenStudy (anonymous):



OpenStudy (anonymous):

this is where the "multiplicity" comes in because the zero at \(-2\) has multiplicity 1, the curve crosses the \(x\) axis at \(-2\)

OpenStudy (anonymous):

that is because \(1\) is an odd number since \(3\) is odd as well, the curve crosses the \(x\) axis at \(0\)

OpenStudy (anonymous):

but \(2\) is even, so the curve does not cross the \(x\) axis at \(5\) it only touches it there

OpenStudy (anonymous):